What decimal value does the 8-bit binary number 10110111 have if:

a) it is interpreted as an unsigned number?

b) it is on a computer using signed-magnitude representation?

c) it is on a computer using one’s complement representation?

d) it is on a computer using two’s complement representation?

e) it is on a computer using excess-127 representation?
Answer #1

a) it is interpreted as an unsigned number?

20 + 21 + 22 + 24 + 25 + 27 = 183

Answer: 183

b) it is on a computer using signed-magnitude representation?

First bit is for sign (0 for positive and 1 for negative)

20 + 21 + 22 + 24 + 25 = 55

Answer: -55

c) it is on a computer using one’s complement representation?

Flip all bits and put a negative sign

Flipped: 01001000

Sign = 23 + 26 = 72

Answer: -72

d) it is on a computer using two’s complement representation?

Most significant magnitude will be of -ve and the rest is positive

- 27 + 20 + 21 + 22 + 24 + 25 + = -128 + 55 = -73

Answer: -73